OpenStudy (anonymous):

what is an area of a circle if the sircles r=8ft

hartnn (hartnn):

area of circle with radius r is \(\large A =\pi r^2\)

OpenStudy (anonymous):

whats thats thing to the left of the r?

hartnn (hartnn):

its \(\large \pi=3.14... or \: \: 22/7\)

hartnn (hartnn):

so, A = 3.14 *r^2 r=8 A = 3.14 * 8^2 =... ?

OpenStudy (anonymous):

25.12?

hartnn (hartnn):

i get A = 3.14*64 = 201.06 ft^2.

OpenStudy (anonymous):

what if the ratio is 20?

hartnn (hartnn):

ratio or radius ?

OpenStudy (anonymous):

same as the last one area andc the r=20

hartnn (hartnn):

r=20 so A = 3.14 *20^2 = 3.14 *400 =... ?

OpenStudy (anonymous):

1256

hartnn (hartnn):

thats correct area.

OpenStudy (anonymous):

hmm well and if d=12?

hartnn (hartnn):

d is the diameter which is half of radius so, r = d/2 = 12/2 = 6 now again same formula for area A =3.14 *6^2 = ... ?

OpenStudy (anonymous):

113

hartnn (hartnn):

that is also correct :)

OpenStudy (anonymous):

and if d=100

hartnn (hartnn):

how would you find r ?

OpenStudy (anonymous):

3.14* it

OpenStudy (anonymous):

so itd be 31400

hartnn (hartnn):

from d which is the diameter , r is half of d so, r= d/2 =100/2 =50 now calculate A.

OpenStudy (anonymous):

huh? :O

hartnn (hartnn):

whenever d = diameter is given, to get radius, divide it by 2.

OpenStudy (anonymous):

so itd be 6?

hartnn (hartnn):

6 ? if d = 100, r= 50

OpenStudy (anonymous):

dikkk what do i do?

hartnn (hartnn):

d= 100, right ? then r=50 A = 3.14 * r^2 = 3.14 * 50* 50 =... ?

OpenStudy (anonymous):

452.16

hartnn (hartnn):

i get something else.

OpenStudy (anonymous):

well thats when you do 3.14*12^2 and then when i did the other part i got 7850

hartnn (hartnn):

i too got 7850

OpenStudy (anonymous):

it was wrong

hartnn (hartnn):

for d=100, A comes out to be 7850...do u want accurate answers ?

hartnn (hartnn):

7853.98

OpenStudy (anonymous):

noo it wasnt on the tes \t so i had to click none of the above it was wrong ):

hartnn (hartnn):

did u understand how to solve such problems ?

OpenStudy (anonymous):

yea.

hartnn (hartnn):

good, that is more important :) that you understand.... any more problems ?

OpenStudy (anonymous):

im dojng a tutorial atm but theyll be more i have till januray 18th to finish up this school year. and in math im 62 percent so i have a lot of tests

hartnn (hartnn):

well, best of luck :)